OpenStudy (anonymous):

The sides of an isosceles triangle are 27 , 27 , and 24yd long. What is the area of the triangle? Do not round any intermediate computations, and round your answer to the nearest tenth.

OpenStudy (anonymous):

how does this one go?

OpenStudy (haichi):

324

OpenStudy (anonymous):

nvrmnd

OpenStudy (anonymous):

Haichi its wrong!

OpenStudy (anonymous):

yea idk lol sorry

OpenStudy (campbell_st):

well you can find the perpendicular... which bisects the base  find the value of h, uby using pythagoras' theorem. then you can find the area

OpenStudy (anonymous):

Thank you campbell_st but I still dont get it by the whole bisects thingy!

OpenStudy (anonymous):

yea thats confusing lol

OpenStudy (anonymous):

Ikr

OpenStudy (anonymous):

Do the pic I think I can help

OpenStudy (campbell_st):

well in an isosceles triangle... the perpendicular from the base... bisects the angle at the top and the base as well... so you end up with a right triangle  and you need the perpendicular height to find the area.

OpenStudy (anonymous):

So how do I find the perpendicular height?

OpenStudy (anonymous):

just kidding nevermind haha

OpenStudy (campbell_st):

use pythagoras' theorem \[27^2 = 12^2 + h^2\] which becomes \[h^2 = 27^2 - 12^2\]

OpenStudy (anonymous):

Thank you so much! I seriously love you !

OpenStudy (campbell_st):

think about this.... you can fold an isosceles triangle exactly in half... and you get 2 identical triangles... the fold line is the perpendicular height... and both triangles are right angled.

OpenStudy (anonymous):

I am so sorry but Im really confused now!

OpenStudy (anonymous):

fr

OpenStudy (campbell_st):

using pythagoras you get \[h^2 = 585\] or \[h = 21.2~ yds\] then the area is \[A = \frac{1}{2} \times 24 \times 21.2 \] hope it helps

OpenStudy (anonymous):

Yes exactly I needed a pic what pic did you use?